OpenStudy (anonymous):

How do I use benchmark estimate for 0.427 and 0.711

OpenStudy (loser66):

\(0.427=\dfrac {427}{1000}\) \(0.711= \dfrac{711}{1000}\) So that you need compare the two using "benchmark" method. That is you need a middle term, let's take 1/2 that is \(\dfrac{500}{1000}\) Now combine them all, we know 427 <500 and 500< 711 so that \(\dfrac{427}{1000}<\dfrac{500}{1000}< \dfrac{711}{1000}\) Hence 0.427 < 0.711 Therefore, your benchmark is 1/2

OpenStudy (loser66):

You can use 0.5 directly to make the comparison. But I am pretty sure your professor wants you to choose either \(1,\dfrac{1}{4},\dfrac{1}{2}\) as candidate for benchmark.